### v1

# Plan Your Travel and Travel with Your Plan: Wide-Horizon Planning and Evaluation via LLM

Abstract
Travel planning is a complex task requiring the integration of diverse real-world information and user preferences. While LLMs show promise, existing methods with long-horizon thinking struggle with handling multifaceted constraints and preferences in the context, leading to suboptimal itineraries. We formulate this as an $L^{3}$ planning problem, emphasizing long context, long instruction, and long output. To tackle this, we introduce Multiple Aspects of Planning (MAoP), enabling LLMs to conduct wide-horizon thinking to solve complex planning problems. Instead of direct planning, MAoP leverages the strategist to conduct pre-planning from various aspects and provide the planning blueprint for planning models, enabling strong inference-time scalability for better performance. In addition, current benchmarks overlook travel’s dynamic nature, where past events impact subsequent journeys, failing to reflect real-world feasibility. To address this, we propose Travel-Sim, an agent-based benchmark assessing plans via real-world travel simulation. This work advances LLM capabilities in complex planning and offers novel insights for evaluating sophisticated scenarios through agent-based simulation.
1 Introduction
Planning a trip xie2024travelplanner ; delarosa2024trippaltravelplanningguarantees is not a straightforward task even for humans, as it requires gathering up-to-date information from various sources, such as hotels, transportation, and points of interest (POI), and integrating this information with preferences and constraints to create a feasible itinerary. With the rapid development of Large Language Models (LLMs) brown2020languagemodelsfewshotlearners ; dubey2024llama ; achiam2023gpt , LLM agents zhu2023ghostminecraftgenerallycapable ; qin2023toolllmfacilitatinglargelanguage greatly improve their capabilities in planning, offering promising opportunities to address the intricate challenges of real-world travel planning.
Existing studies in travel planning chen2024travelagent ; tang2024synergizing ; hao2024largelanguagemodelssolve focus on simple commonsense constraints and coarse-grained user requirements, employing CoT reasoning to integrate collected information into actionable plans. However, they struggle to deliver compelling and viable plans due to their inability to effectively incorporate multifaceted real-world information and implicit constraints into a holistic planning process. For example, a travel plan that is feasible by satisfying the spatial constraints may turn into an unpleasant journey due to ignoring various other aspects, including the special preferences, stamina limitations, or repetitive routes.
Therefore, a well-crafted travel plan necessitates the incorporation of a broader range of information, which in turn requires LLMs to possess enhanced retrieval capabilities and the ability to perform complex planning. We formulate this kind of planning as an $L^{3}$ planning problem with long context (including multifaceted information), long instruction (including CoT and output formatting instruction), and long output (including planning process and a final result). Compared to logical problems, e.g., mathematical reasoning, that demand multi-step analytical processes with strict deductive coherence, $L^{3}$ planning problems require LLMs to simultaneously integrate multifaceted constraints and heterogeneous information, necessitating wide-horizon thinking from various aspects rather than long-horizon thinking. We conduct preliminary experiments to investigate the LLM zero-shot capability of solving the $L^{3}$ planning problem. We implement a naive wide-horizon thinking with aspect-aware question decomposition. Our results reveal that LLMs struggle with addressing $L^{3}$ planning problems with long-horizon thinking, compared to naive wide-horizon thinking.
From the preliminary experiments, we also observe naive wide-horizon thinking exhibits limitations: 1) lack of inter-aspect associations; 2) dependence on artifactual guidance; 3) poor inference-time scalability. To address these limitations, we propose "Multiple Aspects of Planning" (MAoP). In MAoP, instead of directly planning, we introduce a strategist to conduct the pre-planning using a two-stage strategy. The strategist first analyzes the long context and decompose the planning request into various aspects to be considered. For each aspect, the strategist provides short guidance for the planning model to further analyze subsequently. In the second routing stage, the strategist integrates the independent aspects into a coherent planning blueprint, leading to better inference-time scaling performance with more aspects considered. In the actual planning of MAoP, the planning model concentrates on guidance of a single aspect per dialogue turn by following the pre-planning blueprint, progressively constructing a comprehensive planning process via multiple turns of dialogue.
Figure 1: The comparison between long-horizon and wide-horizon thinking reveals distinct cognitive approaches. While long-horizon thinking involves deep exploration of a single reasoning trajectory, wide-horizon thinking incorporates heterogeneous information and constraints from long contexts by considering various aspects. It necessitates parallel consideration of multiple dimensions, which are subsequently integrated to generate comprehensive outputs.
Although MAoP enhances the travel planning capability, the evaluation remains challenging: 1) travel is highly personalized and significantly influenced by real-world dynamics; 2) travel is a consistent process, where previous events can influence the remaining journey. Previous studies xie2024travelplanner ; chen2024travelagent ; shao2024chinatravelrealworldbenchmarklanguage introduce constraint pass rates as metrics, which only reflect the partial feasibility of the plan and the satisfaction of coarse-grained user requirements. These metrics ignore the real-world influence and consistency in real travel scenarios, poorly reflecting actual feasibility and attractiveness.
As the proof is in the pudding, we propose a novel agent-based evaluation framework to simulate a trip based on the real-world environment. A “traveler”, which can be either a group or an individual, functions as an LLM agent to travel by referring to the plan. We design an event-driven sandbox environment integrated with real-world information. For simple events like transiting to a new place, the traveler can leverage map API data to determine the optimal mode of transportation. To simulate complex events like sightseeing, we design an event agent to simulate the experience by integrating up-to-date travel blog posts. Besides considerations of the environment, we meticulously design the traveler agent to make it more aligned with real-world individuals. We design a variety of travelers with distinct features and a traveler stamina engine that simulates stamina exertion. Beyond the "objective" metrics (constraint pass rates), we introduce "subjective" metrics that the traveler evaluates the plan based on the travel experience. The traveler agents offer multi-granularity feedback to assess the experience at multiple levels, ensuring comprehensive and reliable evaluation.
Our contributions can be concluded as follows:
  • Plan Your Travel: We explore the dilemma of addressing the $L^{3}$ problem with long-horizon thinking and propose MAoP to boost the wide-horizon planning capability.
  • Travel with Your Plan: We propose an agent-based evaluation framework that simulates real-world travel to evaluate the feasibility and appeal of the travel plan, offering novel perspectives for assessing complex scenarios.
2 Preliminaries
2.1 Long-Horizon vs. Wide-Horizon
As shown in Figure 1, previous studies wang2023planandsolvepromptingimprovingzeroshot ; yao2023treethoughtsdeliberateproblem ; yao2022react have primarily centered on developing CoT methods and their variants to solve mathematical problems requiring deep exploration along a single trajectory. Deepseek-R1 guo2025deepseekr1 further enhances the long-horizon reasoning ability by leveraging the Reinforcement Learning. Compared to mathematical problems that require long-horizon thinking, we formulate our travel planning problem as an $L^{3}$ planning problem, characterized by long context comprehension, long instruction processing, and long planning process. $L^{3}$ planning problem necessitates LLMs to: 1) extract pertinent information from long contexts (e.g., multiple tour guides, spatial information, user information, and more in Appendix 6); 2) conduct deliberate thinking over multi-dimensional constraints (e.g., real-world constraints and user preferences). $L^{3}$ planning problem does not require the model to reason deeply, as solving mathematical problems with CoTs, but necessitates considering multiple aspects simultaneously in a wide-horizon view. In this paper, we investigate the potential of conducting wide-horizon thinking to solve the $L^{3}$ planning problems.
Table 1: The comparison between long-horizon thinking and wide-horizon thinking in travel planning by evaluating the Feasibility (FEA) and Personalization (PER) scores in Travel-Sim.
Method FEA PER
Qwen 2.5-32B yang2024qwen25
   w/ Artificial Guidance (Long) 23.3 36.2
   w/ Artificial Guidance (Wide) 31.9 44.1
   w/ Self-Gen. Guidance (Wide) 29.4 42.0
DeepSeek-R1 guo2025deepseekr1
   w/ Artificial Guidance (Long) 52.6 62.5
   w/ Artificial Guidance (Wide) 58.9 68.0
   w/ Self-Gen. Guidance (Wide) 55.2 64.1
2.2 Wide-Horizon Thinking with Aspect-Aware Guidance
To explore the influence of simultaneously considering various aspects, we conduct a preliminary experiment to examine the naive wide-horizon thinking in travel planning. We craft guidance (instruction) that includes a series of aspects to be considered when generating a travel plan. To conduct long-horizon thinking (zero-shot CoT) as the baseline, we combine the long context (user information, travel blog posts, spatial information, etc) with the artificial guidance as the input. For naive wide-horizon thinking, we decompose the original guidance into various aspects [^1] Different from the studies perez2020unsupervisedquestiondecompositionquestion ; min-etal-2019-multi that decompose the question into simpler sub-questions to solve it step by step, we decompose the question into aspects to be considered concurrently.. LLMs independently follow corresponding guidance to analyze from different aspects and integrate the whole analysis to produce the output. In Table 1, we find that naive wide-horizon thinking with aspect-aware guidance significantly improves the travel plan quality with better feasibility and personalization scores. Although the wide-horizon thinking demonstrates significant potential, well-designed artificial guidance can further enhance the performance over self-generated guidance. More details can be checked in Section 5 and Appendix D.
3 Plan Your Travel: Wide-Horizon Thinking on Real-world Planning
Based on the preliminaries, we identify three key deficiencies to be improved to enhance the naive wide-horizon thinking: 1) Although aspect-aware thinking is concurrent, its independence inherently limits the capture of inter-aspect associations; 2) the artificial guidance demonstrates superiority, yet its deliberate design remains challenging to implement; 3) as the number of aspects increases, the performance can not be scaled with more inference-time compute. Therefore, we propose the Multiple Aspects of Planning (MAoP), a method that leverages a strategist to conduct the pre-planning to address these deficiencies in naive wide-horizon thinking.
Figure 2: The overview of the MAoP training and inference process.
3.1 MAoP Training
The training of MAoP consists of three stages: reward model training, rejection sampling finetuning for the strategist model, and RL training for the planning model. As shown in Figure 2, we mainly focus on the last two stages and provide more details in Appendix E.
Rejection Sampling Finetuning (RFT) for Strategist Model
In MAoP, as shown in Figure 2, the strategist mainly does two things in pre-planning: 1) decomposing the original instruction and generating the aspect-aware guidance; 2) routing the planning trajectories to capture inter-aspect associations. These core operations solve the mentioned deficiencies in naive wide-horizon thinking. To train such a strategist model, we implement RFT while keeping the planning model frozen  [^2] We do not further implement RL on the strategist model. The RL pipeline has to go through a frozen planning model to get rewards, making it hard to optimize.. For each long context and question, we prompt the strategist model to first conduct pre-planning for $N$ times, and subsequently the planning model to generate final plans. The rejection strategy is to reject the trajectory where all $N$ plans receive reward scores below the threshold.
RL Training for Planning Model
Before RL training, similar to training the strategist model, we first train the planning model with RFT using the same rejection strategy. After RFT (cold start), we conduct online RL training on the planning model to further improve aspect-aware thinking ability. We use GRPO shao2024deepseekmathpushinglimitsmathematical with two kinds of outcome rewards. The first reward comes from the same reward model in the RFT, and the second one is an additional travel plan format reward. We use the KL penalty to prevent severe biases. For loss calculation, we mask the aspect-aware guidance from the strategists, preventing the planning models from generating the guidance by themselves.
3.2 MAoT Inference
3.2.1 Pre-Planning
Decomposition
As shown in Figure 2, the strategist model accepts the long context and question as the input and decomposes the planning request into various aspects. For each aspect, the strategist generates concise guidance to instruct the planning model how to conduct thorough analysis. By parallel sampling the strategist multiple times, we can derive a large amount of aspect-guidance pairs.
Routing
Different from the naive wide-horizon thinking that treats the aspects independently, the strategist additionally selects and aggregates the aspects to route the best planning blueprint. As the experiments in Appendix D.2, we find that if we directly increase the number of aspects by sampling the strategist more times, the planning model only benefits from considering 3 ~5 aspects without further inference-time scaling capability. It is because more aspects introduce more information but also more noise. To address this, the strategist aggregates the number of aspects into a smaller number and constructs a coherent planning blueprint. As shown in Figure 2, subsequent aspect guidance is determined through the influence of preceding multiple aspects. The routing process shifts the burden of considering numerous aspects from the planning model to the strategist model, thus enabling the scalability of considering more aspects to further improve the wide-horizon thinking.
3.2.2 Aspect-Aware Thinking
Based on the planning blueprint constructed by the strategist, the planning model sequentially conducts thinking over the aspects in a way of coherent multi-turn dialogue. With aspect-specific guidance, the planning model can conduct a more focused and in-depth analysis over the long context from this aspect. After multiple turns of profound analysis, the planning model produces the final plan based on the previous wide-horizon thoughts in the last turn of the conversation.
3.3 MAoP Distillation for One-Step Wide-Horizon Thinking
The implementation of MAoP requires two different models and a complicated planning process in multiple turns of dialogue. To accelerate and simplify this process, we directly finetune the models with the MAoP planning synthetic data. Inspired by Deepseek-R1 guo2025deepseekr1 , we especially finetune (SFT) the smaller models like Qwen-2.5 3B yang2024qwen25 , Llama-3.2-Instruct 3B grattafiori2024llama3 , and Deepseek-R1-Distill Qwen-7B guo2025deepseekr1 using 2K synthetic MAoP samples to empower the wide-horizon thinking ability for smaller models. To be specific for data construction, we only preserve the aspect-specific guidance after the routed by the strategist. We compress the entire aspect-aware thinking and the final aggregation into a single turn of conversation as the model output. Therefore, during the inference, the distilled models implement one-step wide-horizon thinking to plan from multiple aspects.
4 Travel with Your Plan: Consistent Evaluation via Agent-based Simulation
Despite the detailed plans we have developed by MAoP, their feasibility in the real world and whether they meet the personalized needs require further verification. As shown in Table 2, previous benchmarks design rule-based/LLM-based metrics, such as pass rates of different constraints. However, travel is a consistent process where previous events influence the remaining journey. By neglecting this consistency, such benchmarks fail to evaluate the actual feasibility and personalized needs in real-world travel scenarios.
As the proof is in the pudding, we design a benchmark using agent-based simulation, dubbed Travel-Sim. In Travel-Sim, the traveler agent experiences a simulated travel referring to the plan and offers feedback from the experience. We use the advanced model, Gemini 2.5-Pro-Exp-0325 google2024gemini , to roleplay the traveler. This simulation mechanism naturally addresses the consistency issue and provides personalized feedback with multiple levels of granularity from the travelers’ perspective.
Table 2: Comparison of evaluation metrics for travel planning between different benchmarks.
Method Rule-based LLM-Judge Multi-Granularity Consistency
TravelPlanner xie2024travelplanner
UnsatChristmas hao2024largelanguagemodelssolve
TravelAgent chen2024travelagent
ITINERA tang2024synergizing
ChinaTravel shao2024chinatravelrealworldbenchmarklanguage
Travel-Sim (ours)
4.1 Consistent Travel Experience Simulation
To simulate a realistic travel experience, we build a sandbox to provide travelers with any real-world information they need, such as time, location, transportation, and even sightseeing experiences from blog posts. Similar to role-playing, we set up detailed profiles for traveler agents, including group sizes, character types, ages, genders, budgets, and preferences. Moreover, we design a stamina engine for the traveler. Stamina is closely related to character types (e.g., young people vs. elderly, family w/ baby vs. family w/o baby) and has a significant impact on the travel experience. The travelers of different types have their own rules for stamina exertion when encountering different events.
Event-Driven State Transition
The traveler follows the plan $p$ to begin the travel within the sandbox. We define the traveler’s state as $c_{n}=\{t,l,s,o,e\}$ at step $n$, where $t$, $l$, $s$, $o$, $e$ represent the current time, location, stamina, outlay, and the ongoing event. The traveler employs a policy $\pi(a_{n}|c_{1},c_{2},\cdots,c_{n-1},p)$, where action $a_{n}\in\mathcal{A}$ and action space $\mathcal{A}$ covers basic actions such as transiting, resting, dining, and sightseeing. Each action the traveler takes leads to a new event, which transits the previous state to the new one. Similar to ReAct yao2022react , the traveler agent first thinks over the current situation from the traveler’s perspective and then makes a decision on the next action.
Real-world Information Integration
When coming to a new city, the traveler usually starts the travel at the train station or airport. The traveler can choose an action, in most cases, the first step is to proceed to the hotel. We utilize a map API to provide various modes of transportation as references. The traveler comprehensively considers stamina (being tired from the long flight), schedule (next event in the plan), and budget to make a decision. For example, if the traveler opts to take the metro to the hotel, the environment updates the traveler’s state with arrival time, new location (hotel), new stamina, etc. Among the events, the simulation of sightseeing events, due to the various experiences and interactions, presents a more significant challenge. Although the traveler agent can not physically visit the POI, we utilize a special sightseeing event agent to generate how the traveler would do the sightseeing by referring to the real experiences in the travel blog posts.
4.2 Multi-granularity Evaluation by Traveler
Evaluation Process
After the trip, the traveler undoubtedly has much to "share". We have the travelers give evaluations on the travel experience from five perspectives (see Appendix F.4 for more details), i.e., experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co). We design a multi-granularity evaluation mechanism from three levels: the traveler reflects on the experience and rates the score after each POI visit, at the day’s conclusion, and after completing the entire journey.
Dataset Construction
We synthesize a variety of traveler profiles, including 16 distinct types of travelers. Each differs in terms of group size, age, gender, stamina level, and preferences. We carefully select 7 Chinese cities that are ideal for tourism as destinations. We have 112 different distinct {traveler, destination, duration} combinations. See more details in Appendix F.3.
5 Experiment
5.1 Evaluation Setup
Methods
For the baseline, we reuse the setup of preliminary experiments, including zero-shot long-horizon thinking and naive wide-horizon thinking. We add an additional baseline that we implement RL to train the models with Long/Artifact. setting using the same training dataset as the MAoP. This additional baseline is to compare the finetuned performance between the long-horizon thinking and the MAoP method. For the MAoP, we train the Qwen 2.5-7B yang2024qwen25 , Qwen 2.5-32B yang2024qwen25 , and Deepseek-R1-Distill Qwen-7B guo2025deepseekr1 as the strategist models and the planning models, respectively. We also include Deepseek-R1 guo2025deepseekr1 and Gemini-2.5-Pro-Exp-03-25 google2024gemini as additional zero-shot planning models. During the MAoP inference, the strategists only route no more than 8 aspects into the planning blueprint for fair comparison.
To simplify MAoP for one-step wide-horizon thinking, we distill the MAoT synthetic data to train Qwen 2.5-3B yang2024qwen25 , Llama 3.2-3B grattafiori2024llama3 , and Deepseek-R1-Distill Qwen-7B. The teachers in the distillation are the Deepseek-R1-Distill Qwen-7B as the strategist model and Gemini 2.5-Pro-Exp-0325 as the planning model. More training details can be checked in Appendix E.
Figure 3: The comprehensive comparison results between MAoP with the baseline methods.
Metrics
We evaluate the MAoP on our Travel-Sim to showcase how MAoP better deals with the $L^{3}$ planning problem via wide-horizon thinking. We introduce four metrics (with details in Appendix G): comprehensiveness (CPH), completeness (CPL), feasibility (FEA), and personalization (PER), where the last two metrics have been used in the preliminary experiment. The first two are rule-based metrics. Comprehensiveness (CPH) evaluates how much relevant information is effectively integrated from the long context into the final plan. Completeness (CPL) evaluates if the travel plan strictly follows the formatting instructions and basic constraints according to the artifactual criteria.
The last two metrics, feasibility (FEA) and personalization (PER), are based on the travel simulation. To evaluate the FEA, also named as Travel Plan Similarity Score (TPSS), we develop an algorithm that calculates the similarity between the trajectory from the travel plan and the trajectory from the simulated travel, as shown in Figure 5 and Algorithm 3. This metric measures the discrepancy between what is planned and the actual execution in the simulation. If the planned trajectory is similar to the simulated one, it means the plan is more feasible. The PER is associated with the feedback of the traveler agent after the simulated travel, indicating if the plan is personalized and suitable for this traveler. The PER is a comprehensive metric, including the evaluation from multiple dimensions and granularities: 1) Multi-dimension: The traveler evaluates the travel experience from five perspectives, i.e., experience (ex), interest (it), arrangement (ar), stamina (st), and cost (co). 2) Multi-granularity: the traveler conducts evaluation from three levels (feedback after visiting every POI, finishing the whole day, and finishing the whole journey). We aggregate the scores from various dimensions and granularities of travel experience, according to Equation 2, to calculate a final PER score.
Table 3: Comparison between MAoP distilled models and MAoP combinations.
Model CPH CPL FEA PER
ex it ar st co agg.
MAoP
   Qwen 2.5-7B (s.) + Qwen 2.5-32B (p.) 64.8 62.5 35.2 59.3 68.4 43.2 49.3 61.5 56.3
   R1-Distill 7B (s.) + R1-Distill 7B (p.) 72.6 76.5 60.7 77.5 86.4 79.3 76.4 87.6 81.4
MAoP Distillation
   Llama 3.2-3B (Distill) 61.3 59.2 52.9 62.0 65.2 63.1 62.5 70.5 65.7
   Qwen 2.5-3B (Distill) 64.2 65.8 53.1 64.2 65.9 63.4 61.4 73.5 66.9
   R1-Distill Qwen-7B (Distill) 78.2 79.2 73.7 84.5 87.2 83.1 76.0 90.2 84.2
5.2 Result and Analysis
5.2.1 Benchmark Performance
As illustrated in Figure 3, for the finetuned models in the first row, the MAoP demonstrates a substantial performance enhancement over the long-horizon thinking (RL w/ Long/Artifact.) baselines, achieving a remarkable 5% to 40% improvement across all planning models. Although trained with the same dataset, MAoP achieves higher scores than RL w/ Long/Artifact. in CPL by better constraint compliance and CPH by integrating more details from the long context to the plan. When compared with naive wide-horizon thinking (Zero-shot w/ Wide/Aritifact.), MAoP excels especially in stronger strategist models, as evidenced by higher FEA and PER scores, suggesting that the strategist plays an important role in the planning process with performance better than artifactual guidance. For advanced models as zero-shot planning models, e.g., Deepseek-R1 and Gemini-2.5, the strategist model also significantly boosts the performance in travel planning compared to zero-shot baselines.
5.2.2 Inference-Time Scaling Capability of the Strategist Model
Figure 4: We experiment two strategist models (Qwen 7B & R1-Distill 7B) and two planning models (Qwen 7B & Gemini) to showcase the scaling capability of the strategist models.
As shown in Figure 4, we find that strategist models can consistently improve the performance by scaling up more considered aspects. We find that even if the model size is the same, the stronger strategist with thinking (R1-Distill 7B) has better scalability, especially for advanced planning models like Gemini 2.5. In contrast, the Qwen 2.5-7B model has limited scalability, because it cannot effectively route a suitable planning blueprint when dealing with increasing aspects.
5.2.3 MAoP Distillation
As shown in Table 3, distilling R1-Distill 7B (s.) + Gemini 2.5-Pro-Exp-0325 (p.) enables 3B sized models to outperform the MAoP combinations even with larger model sizes. For R1-Distill 7B with thinking mode, we also distill the thinking part of planning from the Gemini. Compared to the MAoP combination of the same models, R1-Distill 7B (Distill) achieves better performance. These results demonstrate that distillation from advanced models to smaller models achieves substantial performance improvements, with the enhancement becoming increasingly pronounced as the capability gap between teacher and student models widens, even surpassing the MAoP training models.
5.2.4 Emergence of Spontaneous Behaviors in Consistent Travel Simulation
Instead of strictly obeying the plan, the travel agent refers to the plan and dynamically decides their own itinerary based on various other factors, e.g., stamina, environmental changes, etc. Similar to real-world scenarios, the itinerary is consistent, and subsequent activities are influenced by previous events. Therefore, we observe the emergent behaviors that the traveler agent spontaneously adjusts the predetermined plan, which is not pre-programmed. As the trajectories shown in Figure 5, this elderly couple chooses not to have dinner in Nanmen Lamb Hot Pot because they are too tired to travel to another place, especially since they just suffered from a long train trip to Beijing this morning. In real travel scenarios, when one event goes awry—such as running out of time, excessive fatigue, or attractions being closed—it disrupts the plans for all subsequent activities. This indicates that neglecting consistency issues in travel plan evaluations can lead to significant deviations from real-world scenarios. Travel-Sim surpasses other evaluation benchmarks in assessing consistency by conducting simulations within a real-world sandbox environment.
Figure 5: The example of the trajectories from the generated plan and the trajectory from the simulation. The trajectory of the simulation is not always consistent with the planned one because the traveler agent can change the subsequent itinerary based on the current situation. The FEA score is to calculate the similarity of these two trajectories.
6 Related Work
LLM Planning
The planning capabilities of LLMs have become a key research focus due to their potential as autonomous problem-solving agents wei2023chainofthoughtpromptingelicitsreasoning ; yao2022react ; Wang_2024 . Recent studies have advanced these capabilities through developments in task decomposition, multi-step reasoning yao2023treethoughtsdeliberateproblem ; Besta_2024 , and adaptive planning sun2023adaplanneradaptiveplanningfeedback ; jutrasdubé2024adaptiveplanninggenerativemodels . While these planning algorithms have shown promising results hao2023reasoninglanguagemodelplanning ; hu2024treeplannerefficientcloselooptask , their planning scenarios are limited to simple tasks with a single objective function. For complex real-world planning like travel planning, previous researchers focus on how to preprocess the various information and constraints for LLMs easier to understand but few of them explore how to enhance the planning ability in this complex scenario.
Generative Simulation
LLM agents begin to exhibit strong capabilities in mimicking human behaviors chuang2024demographicsaligningroleplayingllmbased ; wang2023voyageropenendedembodiedagent . Some researchers park2023generativeagentsinteractivesimulacra ; wang2024userbehaviorsimulationlarge ; lin2023agentsimsopensourcesandboxlarge have been investigating the behavioral patterns of human-like LLM agents within sandbox environments, focusing on simulating human social interactions and lifestyles to study their social behaviors. Although simulation-based evaluation has become a prevalent methodology in robotics research chen2024evaluationgenerativeroboticsimulations ; li2024evaluatingrealworldrobotmanipulation , our work represents the comprehensive investigation into evaluating complex task performance using LLM agents within simulated environments.
7 Conclusion
We propose MAoP to enhance wide-horizon thinking for solving $L^{3}$ planning problems, thus addressing the complex challenge of travel planning. In addition, we propose Travel-Sim, an evaluation benchmark that leverages agent-based simulation to offer consistent and multi-granularity evaluation in travel planning scenarios. Our contributions advance the wide-horizon thinking capabilities of LLMs in real-world planning and offer novel insights for evaluating sophisticated scenarios through agent-based simulation.
Acknowledgement
Dongjie Yang and Hai Zhao are with the Department of Computer Science and Engineering, Shanghai Jiao Tong University; Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University; Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3.
This paper was completed during Dongjie Yang’s internship at Xiaohongshu Inc. and was supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400).
Appendix A Limitations and Ethics Statement
A.1 Limitations
While our proposed method (MAoP) and agent-based simulation benchmark (Travel-Sim) demonstrate significant advancements in complex travel planning and evaluation, several limitations warrant discussion.
A.1.1 MAoP Limitations
Though MAoP significantly enhances wide-horizon thinking with deliberate pre-planning by the strategist, MAoP still exhibits a few limitations: 1) MAoP requires the collaboration between the strategist model and planning model, and the conduction of aspect-aware thinking in multiple turns of dialogue. This complicated process costs more inference time than directly CoT. To alleviate this, we propose that the MAoP distillation can compress this complicated process, gaining similar efficiency compared to CoT. 2) We find that weak strategists, especially sizes smaller than 7B, can cause severe performance degradation due to low-quality planning blueprint and poor inference-time scalability. The recommended way to enhance wide-horizon thinking ability for smaller models is to distill the outputs of the advanced model after MAoP training.
A.1.2 Travel-Sim Limitations
The Travel-Sim is the pioneer in evaluating with agent-based simulation, offering a novel solution for complex scenario evaluation, e.g., travel planning. Agent-based simulation and evaluation have several limitations that can be discussed: 1) Similar to other LLM-based evaluation benchmarks, the "traveler", which is an LLM, has the potential bias and unreproducibility over the evaluation. To mitigate this limitation, we stipulate the agent model to be Gemini-2.5-Pro-Exp-0325 [21] using greedy decoding (0 temperature) for reproducibility. We also verify the consistency between LLM-based and human-based evaluation in the Appendix F.5, showcasing the 92% consistency in PER score. 2) Although Travel-Sim takes the consistency and real-world information into account, we do not include all the emergencies that may happen in the real trip, especially some force majeure and human factors. For example, bad weather or delays caused by carelessness also significantly influence the trip. 3) Although Travel-Sim offers consistent and multi-granularity evaluation, the whole evaluation process is time-consuming and costs about 20K tokens per sample. To evaluate the 112 samples of the entire dataset, we cost around \$12 by consuming about 2M tokens in total.
A.2 Ethics Statement
For the traveler profiles in the training dataset and Travel-Sim, we do not use any personal information to collect traveler profiles. The diversity of travelers is manually designed, and data is synthesized using Gemini-2.0-Pro-Exp-0205 [21]. For the sightseeing event agent that uses travel blog posts to simulate travel experience, we collect blog posts from Red Note (Xiaohongshu) that may contain personal information and copyrighted items. Therefore, people using the blog posts should respect the privacy and copyrights of the blog post owner and strictly agree to the license in Appendix B.
Appendix B License
By downloading or using the data, you understand, acknowledge, and agree to all the terms in the following agreement.
ACADEMIC USE ONLY
Any content from the Travel-Sim dataset is available for academic research purposes only. You agree not to reproduce, duplicate, copy, trade, or exploit for any commercial purposes.
NO DISTRIBUTION
Respect the privacy of the personal information of the original source. Without the permission of the copyright owner, you are not allowed to perform any form of broadcasting, modification, or any other similar behavior to the data set content.
RESTRICTION AND LIMITATION OF LIABILITY
In no event shall we be liable for any other damages whatsoever arising out of the use of, or inability to use this dataset and its associated software, even if we have been advised of the possibility of such damages.
DISCLAIMER
You are solely responsible for legal liability arising from your improper use of the dataset content. We reserve the right to terminate your access to the dataset at any time. You should delete the Travel-Sim dataset if required.
You must comply with all terms and conditions of these original licenses, including but not limited to the Google Gemini Terms of Use, the Copyright Rules & Policies of Red Note (Xiaohongshu). This project does not impose any additional constraints beyond those stipulated in the original licenses.
Appendix C Preprocessing Framework
In this section, we introduce how to collect and preprocess the heterogeneous information as the "long context" for travel planning. This context information serves as the information source input both in the training dataset and Travel-Sim.
We break the information and constraints necessary for planning into different categories shown in Figure 6. We leverage our framework to collect and preprocess them for further planning. The framework consists of multiple modules that collaborate with each other.
Figure 6: The decomposition of multifaceted constraints and heterogeneous information to be considered when creating a travel plan.
C.1 Proactive Consultant Agent for Implicit Preferences Elicitation
Users often provide insufficient information, making it challenging for the assistant to devise an effective plan. To address this, we not only extract preferences from user requests but also proactively ask questions to elicit implicit preferences. We design two agents to gather users’ travel and accommodation preferences, respectively. The travel consultant agent collects basic trip details (group size, duration, traveler type) and integrates insights from travel blogs to identify potential challenges for travelers. The accommodation consultant first inquires about hotel preferences and then searches travel blogs and maps to identify optimal hotels. Through iterative interactions, agents can uncover implicit preferences for travel planning and accommodation.
C.1.1 Travel Consultant Agent
We develop an LLM-based agent to serve as a travel consultant, designed to comprehensively dig out the user’s implicit preferences. We design six aspects that the travel consultant agent needs to take into account, as shown in Table 4.
Table 4: Considerations for implicit preferences elicitation.
Type Description
Duration of Visit How long does the user wish to tour, is it one day, two days, or longer?
Group Composition How many people are in the group, what is their approximate age, and are there any special groups, such as the elderly or children?
Attraction Type What type of attractions does the user hope to visit, cultural relics, natural scenery, or shopping and entertainment?
Number of Attractions Does the user want to visit as many attractions as possible, or not too many?
Mode of Transportation Under replaceable circumstances, what mode of transportation does the user prefer: walking, cycling, public transport, taxi, or driving?
Special Requirements Does the user have any special needs or preferences, such as particular requirements for shopping or sightseeing?
Based on the initial user request, the travel consultant agent will proactively inquire about several questions concerning six aspects to better understand the traveler’s situation. Upon gathering sufficient information across these six aspects, the agent will synthesize a comprehensive conclusion in the form of a traveler profile. The agent will subsequently present this profile to the user for review, allowing them to verify the details and provide any necessary additions or corrections through further interaction with the agent.
Instead of asking the same fixed questions, travel consultant agents have the following advantages: 1) Customized Service: we observe that the LLMs will adjust the questions based on the destination and the traveler type, and also give relevant advice to guide the user. 2) Feedback: The agent can receive feedback from the user and raise further interaction to refine the traveler profile.
C.1.2 Hotel Consultant Agent
The selection of the hotel is the key part of travel planning, which seldom previous researchers take into account. The hotel influences not only the accommodation experience but also the planning and logistics of the entire trip, concerning the hotel locations. We develop an LLM-based agent to serve as a hotel consultant agent to recommend the hotel based on travelers’ preferences.
We decompose this process into five phases: 1) Inquiry phase: similar to the travel consultant agent, the hotel consultant agent has to inquire about the user preferences. 2) Online searching phase: after understanding the hotel and POI preferences, the agent can search the internet to find the compliant hotels from the latest tour guides. 3) Map searching phase: the agent leverages the map API to search for hotels near the most-visited locations and conclude a list that meets the requirements. 4) Conclusion phase: the agent recommends the hotels from the travel blog posts and the map API. 5) Feedback phase: the agent receives the feedback and revises the recommendation.
Table 5: Considerations for accommodation planning.
Type Description
Hotel Consistency Does the user prefer to stay in the same hotel throughout the trip, or change hotels according to the attractions?
Accommodation Budget What is the user’s budget for accommodation?
Location Preference Does the user have any requirements for the location of the accommodation, such as being in the city center, near attractions, etc.?
Facility Requirements What facilities does the user require from the accommodation?
To be specific, in the inquiry phase, we consider four aspects, as shown in Table 5. After several proactive interactions, the agent will stop asking if it has collected enough user preferences. Besides the hotel preference, the hotel agent can leverage POI preferences and spatial information (POI clusters) from other modules. In the online searching phase, we use the search API provided by the Red Note (Xiaohongshu) app, a popular Chinese app focused on lifestyle content sharing. We use Red Note (Xiaohongshu) search API to search for blog posts related to hotel recommendations. In the map searching phase, we use the Amap (Gaode) Map API, a Chinese map app similar to Google Maps, to search for hotels near the most-visited locations. The probable most-visited locations are predicted by the agent based on the given information. In the conclusion phase, the agent will recommend the hotels provided by the blog posts and the map API by comprehensively taking the locations and preferences into account.
C.2 Preference-aware POI Selection
As we have the user preferences collected by the travel consultant agent, we leverage the information to find the potential POIs and activities that will appeal to users. We first leverage LLMs to generate several search queries based on user preferences and use the Red Note (Xiaohongshu) API to search the relevant blog posts. Based on the preferences, we use LLMs to extract the POIs that match user preferences. We use the Amap (Gaode) map API to obtain the POI meta information. We finally filter and deduplicate them to get a list of candidate POIs.
C.3 Cluster-based Spatial Optimization
To enhance the coherence of travelers’ journeys and minimize unnecessary back-and-forth travel, similar to ITINERA [9], we cluster the POIs based on geographical distances from each other. This way, visitors can follow a well-planned, spatially coherent route for their tours, for example, visiting the POIs in the same cluster in one day, to enjoy a travel experience that is both efficient and pleasant. We use K-means++ to cluster the POIs, where the number of clusters is determined jointly by the number of travel days and the candidate POIs.
After obtaining the POI clusters, we can describe the spatial distribution characteristics of POIs from two perspectives: intra-cluster features and inter-cluster relationships, to enable LLMs to understand the spatial information.
Algorithm 1 Find Shortest Route in the POI Cluster
Input: locations set $L$, start location name $s$, end location name $e$
Output: shortest path $P$, shortest distance $d_{total}$, step distances $\{d_{i}\}_{i=1}^{n-1}$
function FindShortestRoute($L,s,e$)
     Extract other locations excluding $s$ and $e$: $O\leftarrow\{loc\in L\mid loc\notin\{s,e\}\}$
     Let $n\leftarrow|O|$
     if $n<6$ then
         Compute all permutations of $O$: $\Pi\leftarrow\{\pi\mid\pi\text{ is a permutation of }O\}$
         Find shortest path and distance from $\Pi$:
$(P,d_{total})\leftarrow\arg\min_{\pi\in\Pi}\sum_{i=1}^{|\pi|}\text{CalculateDistance}(C[\pi[i]],C[\pi[i+1]])$
     else
         Get initial path by finding nearest neighbors: $P\leftarrow\text{NearestNeighbor}(L,s,e)$
         Optimize path using 2-opt search: $P\leftarrow\text{TwoOptSearch}(P,C,s,e)$
         Calculate shortest distance: $d_{total}\leftarrow\sum_{i=1}^{|P|-1}\text{CalculateDistance}(C[P[i]],C[P[i+1]])$
     end if
     Calculate step distances: $\{d_{i}\}_{i=1}^{|P|-1}\leftarrow\{\text{CalculateDistance}(C[P[i]],C[P[i+1]])\}_{i=1}^{|P|-1}$
     return $(P,d_{total},\{d_{i}\}_{i=1}^{|P|-1})$
end function
C.3.1 Intra-cluster: Finding the Shortest Routes within the Cluster
Since the POIs in the same cluster are relatively close to each other, travelers have the potential to visit them in sequence within a single day. In order to find the shortest route that does not retrace any part of the path, this problem can be formulated as a variant of the Traveling Salesman Problem (TSP). As the classic Traveling Salesman Problem (TSP) is an NP-hard problem in combinatorial optimization, we use brute force to solve clusters with a small number of POIs (fewer than 6), and for those with a larger number, we use the 2-opt algorithm to find an approximate solution. As shown in Algorithm 1, we have the start location $s$ and end location $e$, which are generally set as the hotel. Locations besides $s$ and $e$ are the POIs in the same clusters. We calculate the shortest route that starts at the hotel and ends at the hotel after visiting the POIs.
C.3.2 Inter-cluster: Relative Cluster Locations to the Hotel
The locations of the POI clusters relative to hotels are also key information for LLM to know the spatial distribution of all POIs. We mainly calculate the direction and distance from the cluster center to the hotel, as shown in Algorithm 2. To make directions easier to understand for LLMs, we can simplify directions into eight basic compass points.
Algorithm 2 Calculate Directions and Distances
Input: Start point $S=(\lambda_{s},\phi_{s})$, set of target points $\mathcal{T}=\{(T_{i},C_{i})\}_{i=1}^{n}$ where $C_{i}=(\lambda_{i},\phi_{i})$
Output: Set of tuples $\mathcal{O}$ containing direction $D_{i}$ and distance $d_{i}$ for each target point $T_{i}$
function CalculateDirectionsAndDistances($S,\mathcal{T}$)
     Initialize empty set $\mathcal{O}$
     for all $(T_{i},C_{i})\in\mathcal{T}$ do
         Convert coordinates to radians: $\lambda^{\prime}_{s},\phi^{\prime}_{s},\lambda^{\prime}_{i},\phi^{\prime}_{i}$
         Calculate geodesic distance $d_{i}$ between $S$ and $C_{i}$
         Compute bearing $\theta^{\prime}$:
$\theta^{\prime}=\arctan 2\left(\sin(\lambda^{\prime}_{i}-\lambda^{\prime}_{s})\cdot\cos(\phi^{\prime}_{i}),\cos(\phi^{\prime}_{s})\cdot\sin(\phi^{\prime}_{i})-\sin(\phi^{\prime}_{s})\cdot\cos(\phi^{\prime}_{i})\cdot\cos(\lambda^{\prime}_{i}-\lambda^{\prime}_{s})\right)$
         Normalize bearing $\theta$: $\theta=(\theta^{\prime}\cdot\frac{180}{\pi}+360)\mod 360$
         Map bearing to cardinal direction $D_{i}$: $D_{i}=\text{directions}[\text{round}(\theta/45)\mod 8]$
         Add tuple $(D_{i},d_{i})$ to $\mathcal{O}$
     end for
     return $\mathcal{O}$
end function
C.4 Entire Preprocessing Workflow
Figure 7: The entire workflow of the preprocessing framework.
Workflow
The preprocessing framework consists of these modules above that work together in collaboration rather than operating independently. As shown in Figure 7, we present the entire workflow of the framework. First, as the user raises a request for planning a trip, the travel consultant agent gathers user preferences and information by interacting with the user. Based on the preferences, we search the relevant blog posts and extract the candidate POIs. We perform clustering on the candidate POIs. In the next step, the hotel consultant inquires about the hotel preference and comprehensively considers all the information (user profile, hotel preference, and spatial information) to recommend hotels. After we have the location of the hotels, we conduct intra-cluster and inter-cluster analysis.
Table 6: The composition of the long context input.
Type Source Description
User Profile & Preferences Travel Consultant The information of the traveler/traveler group and corresponding preferences.
Hotel Information Hotel Consultant The location and other information about where the user lives.
POI Blog Posts Blog Post Search The relevant blog posts that provide sightseeing tips for each POI.
Inter-Cluster Information Spatial Preprocessing The shortest routes within the POI cluster.
Inter-Cluster Information Spatial Preprocessing The relative direction and distance from cluster centers to hotels.
Final Input Context for LLMs
After we collect and preprocess the information, we organize it as the context part of the LLM Input. To give an intuitive view, we structure the context in Table 6. It is noted that for every candidate POI, we search for the most relevant blog post that provides the latest sightseeing tips. We can not depend on the intrinsic knowledge of LLMs about the POIs, which is outdated and inaccurate most of the time. The latest blog posts provide most of the information and tips for visiting the POI, for example, the optimal sightseeing routes, which help LLMs to optimize the itinerary and personalize the plan.
Appendix D Preliminary Experiment Details
D.1 Detailed Setup
Naive Wide-Horizon Thinking with Artifactual Guidance
In the preliminary experiment of exploring naive wide-horizon thinking, we craft the artificial guidance that includes a series of aspects to be considered. We manually design five aspects as shown below:
Naive Wide-Horizon Thinking with Self-Generated Guidance
Instead of using the artifactual guidance, we leverage the LLM to first analyze the aspects to be considered and generate the corresponding guidance. This process can be perceived as the zero-shot strategist, as shown below:
D.2 Scaling Capability of Naive Wide-horizon Thinking
We investigate the potential benefit of scaling inference-time computation with more aspects for naive wide-horizon thinking.
Figure 8: When LLMs consider different numbers of aspects, we analyze the attention pattern of the context (excluding the instruction) on different sizes of Qwen-2.5, varying from 7B, 32B, and 72B (from left to right). (Upper row) We present the distribution of attention scores across the context, aggregated from the output tokens. (Lower row) Additionally, we investigate the impact of context length (number of context tokens) by examining the context awareness scores. Our analysis reveals that larger models exhibit an enhanced capability for wide-horizon thinking, enabling them to focus on a broader range of information within longer contexts by considering multiple aspects simultaneously.
Aspect Qwen 2.5-7B Qwen 2.5-32B Qwen 2.5-72B
CPH $\uparrow$ CPL $\uparrow$ FEA $\uparrow$ PER $\uparrow$ CPH CPL FEA PER CPH CPL FEA PER
0 18.2 40.2 15.2 27.8 39.6 51.1 21.7 34.6 50.4 62.5 37.0 44.0
1 20.3 39.5 14.4 28.5 42.8 53.3 25.4 33.3 51.4 60.3 37.2 43.4
2 25.3 36.2 19.0 30.6 47.0 50.1 31.4 38.7 54.9 60.2 39.2 46.0
3 31.5 37.3 16.3 32.6 49.9 50.4 33.0 42.3 56.5 62.9 41.5 49.7
4 29.8 37.8 14.5 29.4 52.9 47.2 32.0 43.5 60.0 61.7 41.8 52.5
5 28.3 38.0 17.8 27.7 52.7 48.8 31.9 43.1 60.3 59.5 42.1 53.5
Table 7: Comparing the effectiveness of wide-horizon thinking when considering a different number of aspects by quantitative metrics.
Experiment Setup
We investigate scenarios ranging from 0 aspect (i.e., zero-shot CoT baseline without artifactual guidance) to a maximum of 5 artifactual aspects considered. We evaluate how effectively context is considered in the thinking process by analyzing attention patterns with respect to context. Besides attention patterns, we evaluate if wide-horizon thinking is effective and enhances the quality of travel plans by leveraging four metrics: CPL, CPH, FEA, and PER scores. We conduct the experiments on Qwen-2.5 [16] with different sizes, including 7B, 32B, and 72B.
Result Analysis
In the upper part of Figure 8, we show curves of the attention score distributions over the context (excluding the instruction). The attention scores are aggregated from the output tokens and then averaged across layers and attention heads. We observe that larger LLMs (34B and 72B) exhibit higher attention scores across the context when considering more aspects, indicating a strong ability to focus on relevant details. In contrast, the smaller LLM (7B) does not follow this pattern, achieving its highest scores when two aspects are considered. In addition, we introduce another indicator, dubbed context aware score $\mathcal{S}_{c}$, to indicate the proportion of the most attended tokens in the context out of the total, as calculated in Equation 1:
$\begin{gathered}\mathcal{S}_{c}=\frac{1}{n}\sum_{i=0}^{n}\mathbb{I}_{\mathcal{A}^{c}\in\mathrm{Top}k}(\mathcal{A}^{c}_{i}),\end{gathered}$ (1)
where $A^{c}$ denotes the attention scores over the context from the output tokens. In the lower part of Figure 8, for larger models like 72B, as the context length increases, the $\mathcal{S}_{c}$ curves of longer length rise more steeply. It indicates that by considering more aspects, LLMs can capture finer details in longer contexts. This phenomenon emerges only when models are large enough, as smaller LLMs do not exhibit explicit improvements in $\mathcal{S}_{c}$ with long contexts.
As shown in Table 7, larger LLMs demonstrate enhanced travel planning capabilities as they consider more aspects, achieving progressively higher FEA and PER scores. However, this improved performance comes at the cost of formatting compliance, as evidenced by decreasing CPL scores.
Appendix E MAoP Experiment Details
E.1 Training Details
E.1.1 Reward Model
Although we can leverage the results of simulation-based evaluation as the reward signals for RFT and RL training, for the sake of efficiency, we train another reward model. To be specific, we first synthesize 1K travel planning requests. We then separately use Gemini-2.0-Pro-Exp-0205 and Qwen 2.5-32B to directly generate travel plans. For about 2K generated plans, we conduct the simulation-based evaluation to get the PER scores as the reward score labels.
We conduct the pointwise generative reward modeling. We use the travel plan as the input. Instead of directly using the aggregated scalar score calculated in Equation 2, we extract the assessment text of the traveler and corresponding PER score after the whole journey as the output. We use 8 H800 GPUs to finetune the Qwen 2.5-7B model for 5 epochs with a learning rate of 2e-5, global batch size of 64.
E.1.2 Strategist Model
To train a strategist model with RFT, we construct the training dataset consisting of 50K samples with 120 distinct traveler types and 20 Chinese cities. We first conduct RFT on Deepseek-R1-Distill Qwen-7B to get the initial RFT data. We reject the sample that gets the average PER score lower than 40 after sampling $N=3$ times. We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 3 epochs with a learning rate of 2e-5, global batch size of 128.
After training the Deepseek-R1-Distill Qwen-7B, we finally collect 22K RFT data with rejection sampling outputs. We reuse this 22K data to finetune the Qwen 2.5 7B and 32B. We use 8 H800 GPUs to train Qwen 2.5 7B for 3 epochs with a learning rate of 2e-5, global batch size of 64. We use 32 H800 GPUs to train the Qwen 2.5-32B strategist model for 3 epochs with a learning rate of 1e-5, global batch size of 128.
E.1.3 Planning Model
We first train the planning model using the 22K data above as the cold start for subsequent RL training. To train the planning model with RL, we additionally construct the training dataset consisting of 20K samples. We conduct the standard GRPO [19] using the veRL [36] framework. We use 32 H800 GPUs to train Deepseek-R1-Distill Qwen-7B, Qwen 2.5-32B, and Qwen 2.5-7B, for 200 steps with learning rates of 1e-6, 5e-7, 1e-6, respectively, and the KL loss coefficient of 0.001.
E.1.4 Distillation
We reuse the input of the 22K data and use Deepseek-R1-Distill Qwen-7B (s.) + Gemini-2.5-Pro-Exp-0325 (p.) to generate the output. For each sample, we sample the output 3 times. We use the reward model to filter out the output whose reward score is lower than 80. We use the filtered dataset consisting of 15K samples to train the Qwen 2.5-3B, Llama 3.2-3B, and Deepseek-R1-Distill Qwen-7B. For the first two smaller models, we discard the thinking part. For Deepseek-R1-Distill Qwen-7B, we use the thinking part from Gemini-2.5-Pro-Exp-0325. We use 32 H800 GPUs to train Deepseek-Distill Qwen-7B for 5 epochs with a learning rate of 2e-5, global batch size of 128. We use 8 H800 GPUs to train Qwen 2.5 3B and Llama 3.2 3B for 5 epochs with a learning rate of 2e-5, global batch size of 64.
E.2 Inference Details
E.2.1 Inference for Strategist
Decomposition
The first stage of pre-planning is to decompose the planning request into multiple aspects to be considered and provide guidance for further analysis. Here is the prompt used in the RFT and inference:
Routing
If we apply these aspects for subsequent analysis following the naive wide-horizon paradigm, the planning model’s performance will exhibit severe scalability limitations as the number of aspects increases. Therefore, the strategist model reorganizes these aspects and integrates them into a coherent planning blueprint. Here is the prompt used in the RFT and inference:
E.2.2 Inference for Planning Model
Aspect-Aware Thinking
Based on the planning blueprint, the planning model conducts a more focused and in-depth analysis following the order in the blueprint. The whole process of aspect-aware thinking is implemented through multiple turns of dialogue with shared history. We instruct the planning model to follow the guidance to analyze from one of the aspects to analyze in a turn of dialogue, as shown below:
Formatting Output Plan
To integrate the previous analysis, we design formatting instructions for the model to output in a specified format. You can refer to the case study in Section F.2 to see an example of a plan that meets the formatting requirements.
Appendix F Simulation and Evaluation
F.1 Simulation Process
F.1.1 Action Space $\mathcal{A}$
As we have mentioned, there are four kinds of actions permitted to transit to the next state, that is, transiting, resting, dining, and sightseeing. In this section, we illustrate the details of how the travel agent takes these actions.
After the traveler completes the event of the previous state, we prompt the traveler to make the decision for the next action as follows:
Once the traveler chooses an action, we require the traveler agent to give the output formatted as follows:
Transiting: { "decision": "transit", "departure": "…", "destination": "…", "transport mode": "…", "arrival time": "xx:xx", "remaining stamina": …, "total expense": …, "next planned location": "…" }
Resting: { "decision": "rest", "end time": "xx:xx", "remaining stamina": …, "total expense": …, "next planned location": "…" }
Dining: { "decision": "dine", "end time": "xx:xx", "remaining stamina": …, "total expense": …, "next planned location": "…" }
Sightseeing: { "decision": "sightsee", "end time": "xx:xx", "remaining stamina": …, "total expense": …, "next planned location": "…" }
It is noted that only resting, which means resting in place usually (If you’re heading back to the hotel to rest, choose "transit"), does not need additional processing. The following actions, including transiting, dining, and sightseeing, need to integrate with real-world information.
Transiting
If the traveler chooses to transit to the next location according to the plan, there are several modes of transportation, including walking, cycling, public transportation (combination of walking and bus/metro), and taxis. We use the Amap (Gaode) map API to obtain the travel time and costs of these modes of transportation. The traveler agent needs to comprehensively take into account time, cost, preferences, and stamina to make a choice.
Dining
If the traveler chooses to have a meal, we use the Amap (Gaode) map API to find the restaurant near the current location. We provide information on the quality and cost of restaurants for travelers to choose from, which is relevant to the traveler’s preferences and budget. Dining usually takes 0.5-2 hours based on the restaurant type, and is also counted in the time of the sandbox environment.
Sightseeing
If the traveler chooses to go sightseeing at the current location (usually a POI), we simulate the sightseeing experience of the POI. To be specific, we use the Red Note (Xiaohongshu) API to search for travel blog posts related to POI. Since the blog posts are based on real experiences, we design a sightseeing event agent to simulate the probable experience in advance, taking into account the traveler’s preferences, available time, and stamina as follows:
Although we have simulated an experience of the POI in advance, the traveler can consider whether to actually conduct such an experience, considering the current time, stamina, or other factors.
F.1.2 Traveler Stamina Engine
Basic Design
Many researchers focus on analyzing the preference but neglect the type of traveler. The stamina of different types of travelers is diverse, which significantly influences the number of POIs to be visited in one day. In the travel simulation, we design the rule-based engine to adjust stamina expenditure and recovery, which especially vary for different types of travelers. Besides, each type of traveler possesses an initial stamina value, and the travel group is calculated as a whole.
As shown in Table 8, we showcase some examples from Travel-Sim. For younger people, they have higher initial stamina, consume less energy, and recover faster. For elderly people, they would like to have a more relaxed journey and tend to be exhausted more quickly when sightseeing and transiting.
To enable LLMs to intuitively comprehend what stamina represents, we design a stamina-to-state conversion table that translates stamina values into specific states of the traveler: 1) If stamina is greater than 6.0, the state is "Energetic". 2) If the stamina is greater than or equal to 4.0 and less than 6.0, the state is "Good". 3) If the stamina is greater than or equal to 2.0 and less than 4.0, the state is "Slightly Tired". 4) If the stamina is less than 2.0, the state is "Very Tired".
Spontaneous Behavior
Although we do not explicitly instruct the LLM to be aware of the state of the traveler, we find it interesting that the traveler agent can automatically adjust the next action based on the current state. For example, if the traveler feels tired, the next action can be "resting in the restaurant" or "transiting by taxi instead of by bus to the next place". In some cases, the traveler modifies the itinerary or skips the next POI depending on the current state, which illustrates that stamina significantly influences the execution of the actual itinerary. This highlights that previous studies, failing to account for the varying stamina of travelers in planning, have resulted in travel plans that are less feasible.
Table 8: Stamina rules for different types of travelers (from examples in Travel-Sim).
Type Composition {gender, age} [Initial stamina] Description Stamina Rule
Single {male, 32} [8.5] An energetic 32-year-old male traveler loves to explore the natural scenery and historical culture, enjoys in-depth experience of local life through public transportation and walking, especially in traditional food and shopping for special souvenirs, preferring to avoid commercial attractions. sightseeing/1hr: -1; dining: +1; resting/1hr: +1; transiting/1h: bus/metro+0, taxi+0.5, walking-1, cycling-0.5, no cycling
Couple {male, 65}, {female, 62} [6.5] An elderly couple with a passion for culture and history, preferring a leisurely travel pace. They enjoy visiting temples and savoring local cuisine, all while prioritizing rest and comfort during their journeys. sightseeing/1hr: -1.5; dining: +0.5; resting/1hr: +0.5; transiting/1h: bus/metro-1, taxi+0, walking-1.5
Family {male, 32}, {female, 31}, {male, 7}, {female, 4} [7.0] A family of four consisting of a father who loves natural scenery and technology, a mother who loves shopping and searching for local cuisine, and two active children, they prefer a comfortable and convenient way to travel. The family focuses on parent-child experiences and cultural exploration. sightseeing/1hr: -1; dining: +0.5; resting/1hr: +0.5; transiting/1h: bus/metro-1, taxi+0, walking-1.5, no cycling
Family {male, 45}, {female, 41}, {female, 71} [6.5] A family of 45-year-old men, 41-year-old women, and 71-year-old women who prefer relaxing cultural and historical exploration, enjoying local cuisine and special performances, while focusing on the comfort of travel and quality of family time. sightseeing/1hr: -1.5; dining: +0.5; resting/1hr: +0.5; transiting/1h: bus/metro-1, taxi+0, walking-1.5, no cycling
Group {female, 23}, {female, 24}, {female, 27} [7.5] A group of three young female travelers who love to explore local food and culture, fashion shopping and photo check-in travel, prefer a relaxed and comfortable way, enjoy lively nightlife, and do not like to wake up early. sightseeing/1hr: -1; dining: +0.5; resting/1hr: +1; transiting/1h: bus/metro-0.5, taxi+0.5, walking-1, cycling-1
F.2 Case Study
In this section, we showcase an example to intuitively illustrate how the travel simulation works. First of all, we create a traveler agent based on an example in Travel-Sim, e.g., the elderly couple in Table 8. We load the basic information, including the stamina rule and preferences, of the traveler to the system prompt of the LLM agent.
This elderly couple is going to have a 3-day journey in Beijing. They start their journey by arriving at the Beijing West Railway Station at 10:00 AM. Their initial stamina will be subtracted by 2 ($6.5-2=4.5$) because of the exhaustion of traveling to a new city. The couple can choose one of the four actions to go to the next state. We offer four different routes utilizing various modes of transportation, all provided via the map API:
Considering the travelers’ perspective, they decide to take a taxi to their hotel for their next step, as they feel exhausted from the long journey to Beijing:
After the couple arrives at the hotel, they choose to rest for a while before heading to the restaurant to taste the local cuisine, e.g., Peking duck:
(Skipping the states of transiting and dining) After having a wonderful meal in the restaurant, they plan to visit their first attraction, Shichahai, a historic and scenic area in the heart of Beijing, encompassing three interconnected lakes: Qianhai (Front Sea), Houhai (Rear Sea), and Xihai (West Sea). Known for its well-preserved traditional Siheyuan courtyard houses and Hutongs, Shichahai offers visitors an authentic glimpse into old Beijing culture. To create a simulated sightseeing experience in Shichahai, we utilize the Red Note (Xiaohongshu) API to gather blog posts relevant to visiting this area. Based on the gathered content, we then guide the sightseeing event agent to craft an immersive experience from the perspective of the travelers:
Due to the limited space here, we present only a part of the simulation, but it encompasses most scenarios. This case study showcases this consistent travel simulation is based on the integration of real-world information and comprehensively reflects the quality of the generated plan.
Figure 9: An example in Travel-Sim. The traveler group consists of a couple with one female and one male in their 30s. They have an initial stamina of 7.5 and consume less stamina for walking and cycling compared to elders and families with kids. They have specific and detailed preferences that will greatly influence their decision on the itinerary.
F.3 Travel-Sim Dataset Card
As shown in the example in Figure 9, we construct the Travel-Sim dataset that consists of diverse travelers with detailed preferences. Different from previous travel planning benchmarks that only consider single travelers with limited preferences, Travel-Sim accommodates various types of travelers with diverse group compositions, such as individuals, couples, groups, families, and more. Each traveler type may have distinct preferences regarding activities, accommodation standards, budget constraints, travel pace, and cultural experiences.
To build such a dataset, we annotate the profiles of 16 types of travelers and leverage Deepseek-R1 to expand the details. We select the 7 most-visited cities in China as the destinations and combine them with the travelers to generate 112 {traveler, destination} pairs as the evaluation dataset.
F.4 Multi-granularity Evaluation by Traveler
After experiencing the itinerary, the traveler has much to share regarding it. We implement a multi-granularity evaluation mechanism for the travel experience from three levels: the traveler reflects on the experience and rates the score after each POI visit, at the day’s conclusion, and after completing the entire journey. We first ask the traveler agent to think from the perspective of a traveler by outputting the psychological activities. We then ask the traveler agent to rate the score by inspecting five indicators, e.g., travel experience (ex), scenic spot characteristics (it), sightseeing arrangements (ar), stamina exertion (st), and overall expense (co).
For every end of the visiting the POI, we let the traveler assess the POI visiting experience as follows:
For every end of the whole-day itinerary, we let the traveler assess the travel experience of today as follows:
At the end of the multi-day journey, we let the traveler assess the overall travel experience of the entire trip as follows:
F.5 Human Evaluation to Verify the Effectiveness of Simulation-based Evaluation
Although the simulated travel empirically seems to be effective for evaluating the generated plan, we further implement human evaluation to verify if the simulation-based evaluation is consistent with the human evaluation. To be specific, we provide the evaluation results from the simulated travel for humans to check if the evaluation from the traveler agent is reasonable. For example, given the evaluation of the POI travel experience, humans inspect it based on the map and blog post, subsequently deciding whether to endorse the evaluation. If humans agree with the travel agent’s evaluation, the scores remain the same as those of the travel agent. If humans disagree, they have the option to modify the scores.
Three individuals take part in this experiment. We record the modified scores and the deviation values to calculate the agreement rates. It is noted that the scores here are PER-ex, PER-it, PER-ar, PER-st, and PER-co.
As shown in Table 9, the evaluation of traveler agents has high consistency with the human evaluation, with the agreement rate of 92%. We have identified that the PER-co (travel cost) metric exhibits a relatively higher level of inconsistency. The travel agent often buys too many expensive souvenirs, which leads to going over budget. Additionally, they occasionally make mistakes when calculating total travel expenses, especially when accounting for a group of travelers.
Table 9: We implement human evaluation to verify if the simulation-based evaluation is consistent with humans’.
ex it ar st co PER-agg. Agree. rate
R1-Distill 7B (s.) + R1-Distill 7B (p.) 77.5 86.4 79.3 76.4 87.6 81.4 -
Human 69.2 83.7 73.4 74.2 74.5 75.0 92.1%
Appendix G Metrics
We examine four criteria to evaluate the effectiveness of wide-horizon thinking. Comprehensiveness (CPH) and Completeness (CPL) are rule-based metrics, while Feasibility (FEA) and Personalization (PER) are simulation-based metrics.
G.1 Rule-based Metrics
G.1.1 Comprehensiveness (CPH)
Comprehensiveness (CPH) evaluates how much relevant information is effectively integrated from the long context into the final plan. To elaborate, for each POI in the plan, we first extract the corresponding travel guidance. We calculate the similarity between the POI travel guidance and POI-related blog posts in the context. We encode both texts into embedding and calculate the cosine similarity  [^3] We use the paraphrase-multilingual-mpnet-base-v2 model of the sentence-transformers library in https://sbert.net/docs/sentence_transformer/pretrained_models.html.. We calculate the average similarity across all POIs in the plan as the comprehensive score.
G.1.2 Completeness (CPL)
To create an organized travel plan, we must include several essential elements, such as the timeline, destinations, activities, etc. We require LLMs to generate a travel plan that follows the specific output structure and includes the required elements. To evaluate if the generated plan strictly follows formatting instructions for a thorough itinerary, we inspect four criteria as follows:
  1. 1.
    The origin and destination of the entire journey must be at the specified stations or airports.
  2. 2.
    Each day’s itinerary, excluding the first and last day, will begin and end at the hotel where the traveler is accommodated.
  3. 3.
    The introduction for each point of interest’s sightseeing should be formally structured according to the format specified in the prompt.
  4. 4.
    In the travel itinerary, activities should be scheduled to include meal arrangements for both lunch and dinner.
For criteria 1, 2, and 3, we use the regex expression to extract the keywords and verify whether the generated plan satisfies these criteria. For the last one, we leverage the Deepseek-R1 [15] to verify if the itinerary of each day includes arrangements for lunch and dinner. We evaluate each plan against these four criteria, awarding 25 points for each criterion that is met. Therefore, for each plan, we have a maximum of 100.0 points for CPL by summing all the scores.
G.2 Simulation-based Metrics
Because some criteria, e.g., feasibility and personalization, are hard and less persuasive to be simply measured by rule-based metrics, we deal with them via travel simulation based on the real world.
G.2.1 Feasibility (FEA): Travel Plan Similarity Score
We evaluate the feasibility of the generated plan by ensuring that the plan is realistic and executable within the given constraints, such as time and spatial optimization. As shown in Figure 5, we can perceive the generated travel plan as a trajectory of multiple "time-location" pairs. As the traveler conducts a virtual journey based on the generated travel plan, the traveler also produces a trajectory of "time-location" pairs. We aim to assess the feasibility of the generated plan by evaluating whether the plan aligns with the traveler’s itinerary.
To calculate the similarity between two trajectories of "time-location" pairs, there are a few things to consider:
  1. 1.
    Completeness: The events of two trajectories should match. Missing or extra events should be penalized.
  2. 2.
    Chronological order: The events of two trajectories should happen in the same order.
  3. 3.
    Temporal proximity: The identical events of these two trajectories should occur as closely in time as possible. A penalty shall be imposed that corresponds to the temporal discrepancy between identical events. The magnitude of this penalty increases proportionally with the extent of the time gap.
Based on the considerations above, as shown in Algorithm 3, we design an algorithm for calculating the similarity between two trajectories of "time-location" pairs, dubbed Travel Plan Similarity Score (TPSS). We separately deal with the similarity of time and locations and implement dynamic programming to iteratively calculate the score. It is noted that we use TPSS to calculate the similarity of trajectories of one day. For multi-day journeys, we calculate the final TPSS by averaging the daily scores.
Algorithm 3 Travel Plan Similarity Score Calculation
Input: generated_plan_trajectory $T_{g}$, simulated_travel_trajectory $T_{s}$
Output: similarity score (as percentage)
function CalculatePlanSimilarity($T_{g}$, $T_{s}$)
     $m\leftarrow$ length of $T_{g}$
     $n\leftarrow$ length of $T_{s}$
     Initialize $dp[m+1][n+1]$ with zeros
     for $i\leftarrow 1$ to $m$ do
         for $j\leftarrow 1$ to $n$ do
              $score\leftarrow$ CalculateMatchScore($T_{g}$[i-1], $T_{s}$[j-1])
              $dp[i][j]\leftarrow\max(dp[i-1][j],dp[i][j-1],dp[i-1][j-1]+score)$
         end for
     end for
     $max\_score\leftarrow\min(m,n)$triangleright Ideal maximum score
     $similarity\leftarrow dp[m][n]/max\_score$
     $completeness\_penalty\leftarrow\min(m,n)/\max(m,n)$triangleright Penalize for missing/extra activities
     $final\_similarity\leftarrow similarity*completeness\_penalty*100$
     return $final\_similarity$
end function
function CalculateMatchScore(item1, item2)
     $time\_score\leftarrow$ TimeDiffScore(item1.time, item2.time)
     if item1.location = item2.location then
         $location\_score\leftarrow 1$
     else
         $location\_score\leftarrow 0$
     end if
     return $(time\_score+location\_score)/2$
end function
function TimeDiffScore(time1, time2)
     $diff\leftarrow|$time2 - time1|||| in hours
     return $\max(0,1-diff/2)$triangleright Linear decrease within 2 hours
end function
G.2.2 Personalization (PER): Aggregated Assessment from Travelers
We evaluate the personalization of the generated plan by inspecting if the travel plan meets the unique needs and preferences of the traveler agent in the simulation. As we have mentioned, our multi-granularity evaluation mechanism via traveler’s assessment in Appendix F.4, we further introduce how we aggregate the scores from multi-granularity assessment in Equation 2 as follows:
$\mathcal{S}_{PER}=\alpha_{1}\mathcal{S}_{whole\_travel}+\alpha_{2}\frac{1}{N}\sum_{i}^{N}(\beta\mathcal{S}_{day\_i}+\gamma\mathcal{S}_{POI\_day\_i}).$ (2)
As each POI has a score after sightseeing, we average the scores of each POI to be the overall POI score $\mathcal{S}_{POI\_day\_i}$ in $i^{th}$ day. We weight the $\mathcal{S}_{POI\_day\_i}$ and the score of the whole-day itinerary $\mathcal{S}_{day\_i}$ by $\beta$ and $\gamma$ as the aggregated score of $i^{th}$ day. We set $\beta$ as 0.6 and $\gamma$ as 0.4. We average the aggregated score of each day and weight it with the score of the whole travel by $\alpha_{1}$ and $\alpha_{2}$. We set $\alpha_{1}$ as 0.6 and $\alpha_{2}$ as 0.4. We normalize the $\mathcal{S}_{whole\_travel}$ to 100.0 as the final score.